OpenStudy (anonymous):

You have six $1 bills, eight $5 bills, two $10 bills, and four $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill. What is P($1, then $10)? 77/190 3/100 3/95 2/5

OpenStudy (anonymous):

i'm trying to get this math test done and i need help

OpenStudy (jtvatsim):

alright, what is your thought process on the question so far? how are you thinking to approach the problem? :)

OpenStudy (anonymous):

the probability the first one is a $1 is the number of $1 bills over the total number of bills

OpenStudy (anonymous):

i.e. \(\frac{6}{20}=\frac{3}{10}\)

OpenStudy (anonymous):

the probability the second is a $10 given that the first one is a $1 is \(\frac{2}{19}\) the \(19\) because you have already chosen a $1 bill and not replaced it

OpenStudy (anonymous):

to find your answer, multiply those numbers i.e. compute \[\frac{3}{10}\times \frac{2}{19}\]

OpenStudy (jtvatsim):

yup, that's the way it's done! :)